### v1

@@ 4.5.3.1 Performance Comparison for Informative Task
RoBERTaMFT [57] 88.90 - - 87.40
DMCC [58] 92.24 92.24 92.23 92.23
CrisisSpot 97.58 96.70 99.80 98.23
We have used various notations for evaluation metrics: accuracy as (Acc), precision as (Prec), recall as (Rec), and F1-Score as (F1). All these metrics are presented in percentage (%) in all the Tables of the paper.
4.5.3.2 Performance Comparison for Humanitarian Task
[…]
Table 7: Performance Comparison on CrisisMMD for Humanitarian Task
Model Acc Prec Rec F1
BERT [37] 81.25 82.13 82.72 82.42
ResNet [51] 81.33 81.29 81.98 81.63
Visual BERT [52] 84.05 82.22 83.56 82.88
VilT [54] 83.30 83.82 84.92 84.37
ViLBERT [53] 84.33 84.95 86.52 85.73
SCBD [18] 80.62 80.77 80.52 80.64
MANN [19] 78.42 78.59 78.42 78.50
CAMM [55] 82.19 82.31 81.88 82.09
[…]
RoBERTaMFT [57] 84.90 - - 80.40
DMCC [58] 88.00 87.95 87.80 87.72
CrisisSpot 90.01 90.87 90.04 90.13
It is important to highlight that the humanitarian task entails multi-class classification, posing a challenge due to uneven label distributions. Notably, our proposed approach achieves notable performance improvements over existing models, even without employing data augmentation to balance the dataset. This improvement can be attributed to our incorporation of word-level and sentence-level features. This approach considers both local and global contexts within the modalities, resulting in a richer context that ultimately enhances the overall model performance. Additionally, the incorporation of semantic text features such as emotion features, sentiment features, the Crisis Informative Score, and various social interaction scores further enhances the model’s understanding of the context. In conclusion, these thorough comparisons highlight the effectiveness of CrisisSpot within the realm of disaster management.
4.5.4 Performance Comparison of TSEqD
To validate the effectiveness of CrisisSpot, we have compared it with existing unimodal and multimodal state-of-the-art benchmarks on our TSEqD dataset.
4.5.4.1 Performance Comparison for Informative Task
[…]
Table 8: Performance Comparison on TSEqD for Informative Task
Model Acc Prec Rec F1
BERT [37] 76.42 78.29 79.35 78.81
ResNet [51] 77.85 79.54 80.23 79.89
Visual BERT [52] 78.56 80.21 81.28 80.74
VilT [54] 78.92 79.65 81.28 80.45
ViLBERT [53] 79.99 82.24 86.86 84.48
SCBD [18] 80.79 83.16 89.58 86.25
MANN [19] 80.14 82.76 88.55 85.55
CAMM [55] 81.08 83.89 90.34 86.99
[…]
RoBERTaMFT [57] 78.23 82.55 88.25 85.30
DMCC [58] 81.10 84.15 90.13 87.03
CrisisSpot 84.42 84.57 96.20 90.01
4.5.4.2 Performance Comparison for Humanitarian Task
We have performed a comparative analysis of CrisisSpot, evaluating it against both unimodal and multimodal state-of-the-art methods in the context of Humanitarian Tasks. As depicted in Table 9, our method achieves an accuracy of 80.55%, a precision of 81.87%, a recall of 82.20%, and an F1-score of 82.03% for the Humanitarian Task. Our method outperforms both the unimodal and multimodal state-of-the-art benchmarks by significant margins. Furthermore, we assessed our results in comparison to FixMatchLS (referenced in [56]), RoBERTaMFT (cited in [57]), and DMCC (mentioned in [58]). Our method exhibited performance improvements over FixMatchLS by 8.32% in accuracy, 6.92% in precision, 3.24% in recall, and 4.07% in F1-score. When compared to RoBERTaMFT, CrisisSpot shows significant improvements of 9.32% in accuracy, 7.32% in precision, 3.4% in recall, and 5.67% in F1-score, respectively. Furthermore, in comparison to DMCC, our method demonstrated performance gains of 2.45% in accuracy, 4.72% in precision, 3.4% in recall, and 4.07% in F1-score. Furthermore, we benchmark CrisisSpot against SCBD[18], MANN[19], and CAMM[55], demonstrating our method’s substantial performance improvements across all evaluated metrics compared to these multimodal approaches. These comprehensive comparisons underscore the strong performance of CrisisSpot.
Table 9: Performance Comparison on TSEqD for Humanitarian Task
Model Acc Prec Rec F1
BERT [37] 72.22 74.44 75.84 75.13
ResNet [51] 72.85 75.54 76.23 75.88
Visual BERT [52] 74.56 76.21 78.28 77.23
VilT [54] 74.02 75.23 78.25 76.71
ViLBERT [53] 76.00 78.27 80.19 79.22
SCBD [18] 77.50 78.03 78.58 78.30
MANN [19] 76.80 77.25 78.95 78.09
CAMM [55] 78.00 79.43 80.34 79.88
[…]
RoBERTaMFT [57] 71.23 74.55 78.25 76.36
DMCC [58] 78.10 77.15 78.80 77.96
CrisisSpot 80.55 81.87 82.20 82.03
@@ 4.5.5 Performance Gain Analysis
We assess the performance improvement achieved by CrisisSpot through a systematic analysis. Initially, we evaluate our model’s performance with different combinations of modalities and subsequently explore the impact of the attention mechanism. Subsequently, we evaluated the effect of incorporating different combinations of Social Context Features in our method.
4.5.5.1 Analyzing Combinations of Modalities
We scrutinize the performance of the CrisisSpot using diverse combinations of modalities. Due to the varying nature of the different modalities, it becomes imperative to analyze the effect of various combinations of modalities used in our approach. The inputs to our method encompass three distinct modalities: textual modality (T), visual modality (I), and Social Context Features (SCF). The performance comparisons of these diverse modality combinations are detailed in Table 10.
Table 10: Modality-based Comparison on Informative Task
Model Acc Pre Rec F1
CrisisSpot (w/o Image) 91.72 92.20 94.66 93.41
CrisisSpot (w/o Text) 92.56 93.84 95.35 94.14
@@ 4.5.5.1 Analyzing Combinations of Modalities
‘T’ and ‘I’ represent Text and Image respectively.
CrisisSpot (I+T+SCF) represents our proposed method as shown in Table 10, encompassing all three modalities. To assess the significance of the image input, we exclude it while preserving text and Social Context Features, resulting in the model CrisisSpot (w/o Images). Similarly, the CrisisSpot (w/o Text) model utilizes images and Social Context Features as input, omitting text data. Furthermore, to highlight the importance of Social Context Features, we exclude them from our model. This version is referred to as CrisisSpot (w/o SCF), using text and images as inputs.
The CrisisSpot (w/o SCF) exhibits substantial performance compared to the CrisisSpot (w/o image) with improvements of 2.9% in accuracy, 3.66% in precision, 5.14% in recall, and 5.09% in F1-score, respectively. This improvement can be attributed to the comprehensive consideration of both modalities in the CrisisSpot (w/o SCF). In the CrisisSpot (w/o image), the focus is primarily on textual and Social Context Features, which limits its performance. Additionally, an attention mechanism was not applied in the CrisisSpot (w/o image) due to the presence of only a single modality.
CrisisSpot (T+I+SCF), which combines textual modality, visual modality, and Social Context Features, outperforms all other variants across all evaluation metrics. In comparison, CrisisSpot (T+I+SCF) exhibits a significant performance boost, surpassing CrisisSpot (w/o SCF) by 2.73% in accuracy, 0.84% in precision, 3.38% in recall, and 2.09% in F1-score. Moreover, it outperforms CrisisSpot (w/o Text) by 2.95% in accuracy, 2.84% in precision, 4.45% in recall, and 4.09% in F1-score. This significant enhancement in the CrisisSpot (T+I+SCF) compared to its variants comes from the effective utilization of multimodal information and Social Context Features present in the posts. These results underscore the efficiency of our proposed multimodal approach and emphasize the significance of each modality employed in the model.
4.5.5.2 Analysis of Inference Time
Inference is the stage where a trained model makes predictions or classifications based on input data. Inference time refers to the time taken by the model to make the predictions. We measure inference time in milliseconds (ms).
Table 11: Analysis of Inference Time
Model TIT(ms) PSIT(ms)
BERT [37] 2685 1.75
ResNet [51] 1427 0.93
CAMM [55] 2823 1.84
VilT [54] 3636 2.37
MANN [19] 4495 2.93
DMCC [58] 4541 2.96
SCBD [18] 6903 4.50
[…]
RoBERTaMFT [57] 7517 4.90
ViLBERT [53] 8437 5.50
Visual BERT [52] 9971 6.50
CrisisSpot 6520 4.25
TIT: Total Inference Time, PSIT: Per Sample Inference Time
Number of samples: 1534
[…]
4.5.5.3 Analysis of Attention Mechanism
We conducted a performance analysis of CrisisSpot with and without an attention mechanism. We denote our proposed model with attention as CrisisSpot (With Att) and without attention as CrisisSpot (W/o Att).
CrisisSpot (w/o Att): It refers to the implementation of the proposed model without attention applied to both textual modality and visual modality. Here, we did not apply any kind of attention to the text and image features, resulting in not capturing the complex interactions between textual and visual modalities. The resulting model gets an accuracy of 95.86%, precision of 94.32%, recall of 98.72, and F1-score of 96.47% as shown in Table 12.
Table 12: Attention-based Comparison for Informative Task
Model Acc Prec Rec F1
CrisisSpot (w/o Att) 95.86 94.32 98.72 96.47
CrisisSpot (with Att) 97.58 96.70 99.80 98.23
CrisisSpot (with Att): This description relates to the implementation of our proposed approach with an attention mechanism applied to multimodal content. In this particular variant, attention is applied to the word embeddings obtained from the text encoder based on transformers and the image embeddings from the image encoder. This attention mechanism is utilized to identify both similar and dissimilar patterns within the data. As the CrisisSpot (with Att) models the interaction between the text and image modalities, it achieves an accuracy of 97.58%, precision of 96.70%, recall of 99.80, and F1-score of 98.23%, as shown in Table 12.
4.5.5.4 Analysis of Social Context Features
We demonstrate the significance of the Social Context Features (SCF) employed in the CrisisSpot. The SCF discussed in this section are User Informative Score (UIS), Crisis Informative Score (CIS), and other features. We denote the proposed model without User Informative Score as CrisisSpot (w/o UIS). Similarly, we denote the proposed model without Crisis Informative Score as CrisisSpot (w/o CIS), and the proposed model without other features as CrisisSpot (w/o other features). The other features refer to the Sentiment Quotient, Emotion Quotient, and User Engagement Metrics.
CrisisSpot (w/o UIS): As shown in Table 13, the CrisisSpot (w/o UIS) attains an accuracy of 96.08%, a precision of 95.23%, a recall of 98.31%, and an F1-score of 96.74%. The omission of User Informative Score (UIS) integration in our model has resulted in the neglect of user credibility assessment. Extensive experimentation has demonstrated the pivotal role UIS plays in assessing the informativeness of social media content during disaster events. A high UIS score indicates that a user is contributing more substantiated and informative content.
Table 13: Social Context Feature-based Comparison for Informative Task
Model Acc Prec Rec F1
CrisisSpot(w/o UIS) 96.08 95.23 98.31 96.74
@@ 4.5.5.4 Analysis of Social Context Features
‘U’ represents User Informative Score (UIS)
‘C’ represents the Crisis Informative Score (CIS)
‘OF’ represents the Other Features
CrisisSpot (w/o CIS): As demonstrated in Table 13, the CrisisSpot (w/o CIS) attains an accuracy of 97.03%, a precision of 96.18%, a recall of 99.21%, and an F1-score of 97.67%. The absence of the Crisis Informative Score in our approach implies that we are overlooking the significance of crisis-related vocabulary within a text. The CIS score essentially measures the presence of crisis-related words in a given text. A higher CIS score indicates that the text contains a greater number of crisis-relevant terms, thereby signifying a higher level of informative content related to the crisis.
CrisisSpot (w/o other features): As presented in Table 13, the CrisisSpot (w/o other features) obtains an accuracy of 96.63%, a precision of 95.75%, a recall of 98.85%, and an F1-score of 97.28%. Neglecting to incorporate emotion, sentiment, and user engagement metrics into the model does not accurately capture the emotional context of the content, potentially leading to misinterpretations or oversights in understanding the underlying sentiment. Secondly, the absence of user engagement metrics could hinder the model’s ability to gauge the relevance and impact of content, potentially overlooking important interactions. Overall, without considering these features, the model may perform poorly.
CrisisSpot (CIS + UIS + Other Features): This pertains to the suggested model. As indicated in Table 13, our proposed method attains an accuracy of 97.58%, a precision of 96.70%, a recall of 99.80%, and an F1-score of 98.23%. Our proposed approach demonstrates the most favorable outcomes as it integrates all the Social Context Features, contributing to improved crisis content classification.
4.5.6 Model Component Analysis
In Table 14, we illustrate the significance of each module and how they influence the performance of our proposed model. The various variants are CrisisSpot (w/o SCFM), which denotes the proposed model without a Social Contextual Feature Module (SCFM); CrisisSpot (w/o MGLM), which denotes the proposed model without a Multimodal Graph Learning Module (MGLM); CrisisSpot (w/o FEIM), which denotes the proposed model without the feature extraction and interaction module (FEIM); and CrisisSpot (FEIM + MGLM + SCFM) denotes the proposed model.
Table 14: Model Component Comparison for Informative Task
Model Acc Prec Rec F1
CrisisSpot(w/o SCFM) 94.63 95.86 96.42 96.14

### v2

@@ 5.5.3 Performance Comparison of TSEqD
RoBERTaMFT (Yoon et al., 2023) 78.23 82.55 88.25 85.30 74.23 74.55 78.25 76.36
DMCC (Rezk et al., 2023) 81.10 84.15 90.13 87.03 78.10 77.15 78.80 77.96
CrisisSpot 84.42 84.57 96.20 90.01 80.55 81.87 82.20 82.03
The results, as presented in Table 7, showcase CrisisSpot’s impressive performance metrics. For the Informative Task, CrisisSpot achieves an accuracy of 84.42%, a precision of 84.57%, a recall of 96.20%, and an F1-score of 90.01%. These metrics demonstrate significant improvements over both unimodal and multimodal state-of-the-art benchmarks, highlighting CrisisSpot’s superiority in disaster content classification. In comparison to RoBERTaMFT, CrisisSpot outperforms it by 8.68% in accuracy and 10.83% in the F1-score. Similarly, when benchmarked against DMCC, CrisisSpot exhibits performance gains of 3.32% in accuracy, 0.42% in precision, 6.07% in recall, and 2.98% in F1-score. These results emphasize CrisisSpot’s consistent superiority over other methods across various evaluation metrics. Furthermore, CrisisSpot is compared against SCBD (Abavisani et al., 2020), MANN (Gao et al., 2020), and CAMM (Khattar and Quadri, 2022), all of which represent multimodal state-of-the-art methods. Notably, CrisisSpot outperforms these methods by substantial margins across all evaluation metrics. This comprehensive comparison underscores the robustness and effectiveness of CrisisSpot in the context of disaster content classification. In conclusion, the superior performance exhibited by CrisisSpot reaffirms its position as an advanced solution for disaster content classification. By surpassing existing benchmarks, CrisisSpot demonstrates its ability to leverage both unimodal and multimodal information effectively, thereby enhancing decision-making processes in crisis scenarios.
@@ 5.5.4 Performance Gain Analysis
We assess the performance improvement achieved by CrisisSpot through a systematic analysis on CrisisMMD for Informative tasks. Initially, we evaluate our model’s performance with different combinations of modalities and subsequently explore the impact of various attention mechanisms. We evaluate the effect of incorporating different combinations of model components. Finally, we analyze various language, vision, and graph methods.
Table 8: Performance Gain Analysis of Different Model Variants
Caption Model Acc Pre Rec F1
Modality-based Comparison CrisisSpot (w/o Image) 92.56 93.84 95.35 94.14
CrisisSpot (w/o Text) 91.72 92.20 94.66 93.41
@@
Attention-based Comparison No Attention 95.86 94.32 98.72 96.47
Self Attention 95.95 94.48 98.85 96.58
Cross Attention 96.12 95.26 98.97 97.21
IDEA 97.58 96.70 99.80 98.23
Social Context Feature-based Comparison CrisisSpot(w/o UIS) 96.08 95.23 98.31 96.74
@@
Model Component Comparison CrisisSpot(w/o SCFM) 94.63 95.86 96.42 96.14